OpenStudy (anonymous):

will someone please help me with adding and subtracting fractions please i will give out medals!!!!

OpenStudy (the_fizicx99):

Sure

OpenStudy (anonymous):

What is the difference? -2/15 - (-9/15) @tHe_FiZiCx99

OpenStudy (the_fizicx99):

they have the same denominator lol just simply subtract -2 - (-9)

OpenStudy (anonymous):

okay would you do the next one the same way???

OpenStudy (the_fizicx99):

Depends, the same denominator? Addition and subtraction require same denominator. Multiplication and division don't

OpenStudy (anonymous):

What is the sum? 2/3 + 7/15

OpenStudy (anonymous):

depends.. if they have common denomenators, then yes.

OpenStudy (anonymous):

find common denominator

OpenStudy (the_fizicx99):

LCD.

OpenStudy (the_fizicx99):

Can you find it?

OpenStudy (anonymous):

which would be 3?? right

OpenStudy (anonymous):

??

OpenStudy (the_fizicx99):

No, 15. gtg :p gl

OpenStudy (anonymous):

ok

OpenStudy (lovelyanna):

The sum of 2/3+7/15= 10/15+7/15= 17/15 or 1 2/15